# Learning State-Invariant Representations of Objects from Image Collections with State, Pose, and Viewpoint Changes

### v1

@@ Abstract
We add one more invariance — the state invariance — to the more commonly used other invariances for learning object representations for recognition and retrieval. By state invariance, we mean robust with respect to changes in the structural form of the object, such as when an umbrella is folded, or when an item of clothing is tossed on the floor. Since humans generally have no difficulty in recognizing objects despite such state changes, we are naturally faced with the question of whether it is possible to devise a neural architecture with similar abilities. To that end, we present a novel dataset, ObjectsWithStateChange, that captures state and pose variations in the object images recorded from arbitrary viewpoints. We believe that this dataset will facilitate research in fine-grained object recognition and retrieval of 3D objects that are capable of state changes. The goal of such research would be to train models capable of generating object embeddings that remain invariant to state changes while also staying invariant to transformations induced by changes in viewpoint, pose, illumination, etc. To demonstrate the usefulness of the ObjectsWithStateChange dataset, we also propose a curriculum learning strategy that uses the similarity relationships in the learned embedding space after each epoch to guide the training process. The model learns discriminative features by comparing visually similar objects within and across different categories, encouraging it to differentiate between objects that may be challenging to distinguish due to changes in their state. We believe that this strategy enhances the model’s ability to capture discriminative features for fine-grained tasks that may involve objects with state changes, leading to performance improvements on object-level tasks not only on the new dataset we present, but also on two other challenging multi-view datasets such as ModelNet40 and ObjectPI.
@@
Deep Metric Learning, Object Recognition and Retrieval, Pose-invariant and State-invariant Embeddings, Fine-grained Tasks, Multi-modal Dataset with Multi-view Images of Objects in Different States and Text Descriptions
@@ 1 Introduction
The fact that the appearance of an object depends significantly on its pose vis-a-vis that of the camera has received much attention in the research literature. However, there exists a more confounding cause for major changes in object appearance — the change in the state of the object. Consider, for example, objects like umbrellas that may be open or closed, books on a table that may also be open or closed, furniture items that are likely to be folded up when not in use, items of clothing and linen in different forms depending on how they are put away, etc. Fig. 1 shows some examples of such objects from our ObjectsWithStateChange dataset.
Figure 1: In addition to the pose, many commonly occuring objects also exhibit significant changes in their appearance when their state changes. This figure displays several examples of such objects in our ObjectsWithStateChange dataset.
We posit that with regard to the objects we enounter in our daily lives, it is just as likely as not that you will encounter objects whose appearance depends significantly on what state we find them in. That makes important the following question: Can modern computer vision algorithms effectively recognize objects despite the changes in their state? That is, in addition to achieving pose and viewpoint invariances, is it possible to also achieve invariance with respect to state changes? We believe that the work we present is a step in the direction of answering those questions.
@@
Obviously, any exploration of the answers to the questions posed above must begin with the creation of a dataset that captures not only pose and viewpoint appearance changes, but also state changes. To that end, a primary goal of this paper is the presentation of a new dataset we have created — the ObjectsWithStateChange dataset (hrta. OWSC). We believe that this dataset nicely augments the other well-known datasets — ModelNet40, ObjectPI, and FG3D — that are only meant to train and test for invariances with respect to pose and viewpoint. We believe that the OWSC dataset would facilitate creating image representations that would also be invariant to highly nonlinear shape deformations caused by state changes in addition to being invariant to changes in viewpoint, pose, illumination, etc.
To demonstrate the usefulness of the OWSC dataset, we also propose an extension to the dual-encoder model presented in [11] that learns object and category representations in separate embedding spaces simultaneously from multiple images of an object under different transformations. The model learns discriminative features by comparing images of a pair of objects that are randomly sampled from the same category. In order to accommodate the changes in object state, we extend [11] by proposing a curriculum learning strategy that uses the inter-object distances in the learned embedding space after each epoch to guide the sampling process. Specifically, this involves sampling images of similar objects from the same category (by retrieving neighboring objects from the same category, with and without state changes, in the learned embedding space) and from diverse categories (by dividing the learned embedding space into finer partitions as training progresses and sampling neighboring objects from the same partition). By doing so, we encourage the model to differentiate between harder-to-distinguish objects, enhancing its ability to capture discriminative features for fine-grained tasks in the presence of changes in state.
@@
  • We introduce a new dataset called ObjectsWithStateChange that incorporates state changes in object appearance that are in addition to the more commonly used transformations related to changes in pose and viewpoint. We evaluate several state-of-the-art approaches on eight different category and object-level classification and retrieval tasks using this dataset.
  • We propose a new mining strategy to learn discriminative object-level representations more effectively that appears to work well even in the presence of state changes for the objects. This strategy leads to performance improvements on object-level tasks not only on the proposed ObjectsWithStateChange dataset but also on two other publicly available multi-view datasets, ObjectPI and ModelNet-40.
@@ 2 Related Work
(A) Multi-view Object Datasets: Prior work has introduced several multi-view object datasets for recognition and retrieval tasks. Datasets such as ModelNet-40 [1], and FG3D [2] comprise multi-view images of CAD models of synthetic objects. These datasets are intended for learning shape of objects and lack color and texture information, that we see in real-world objects. Datasets such as ETH-80 [3], W-RGBD [4], RPC [9], MIRO [12], MVP-N [10] capture real-world objects placed in a controlled environment with uniform illumination and a plain background while the ObjectPI [7] dataset captures objects in the wild with complex backgrounds. In all of these datasets, multi-view data is captured of objects placed in a particular pose from pre-defined viewpoints. ObjectNet [8] is a test dataset meant for evaluation of category recognition in a real-world setting with arbitrary object backgrounds, rotations, and viewpoints. Toybox [6] is a small video dataset capturing basic transformations such as rotation, translation, and zoom.
@@
(B) Deep Metric Learning Losses and Exemplar Mining:
Ranking Losses:
In deep metric learning, ranking losses [13, 14, 15] aim to optimize relative pairwise distances among exemplars, pulling similar samples closer together while pushing dissimilar ones apart. Selecting informative exemplars [16, 17, 18, 19, 20, 21, 22, 23] is crucial but adds computational costs and memory. To reduce the training complexity proxy-based methods [24] define a proxy embedding for each class and optimize the sample-to-proxy distances, which however does not capture the rich sample-to-sample relations that are particularly crucial for fine-grained tasks.
@@
Curriculum Learning: As introduced in the seminal work [25], it organizes learning tasks by gradually increasing complexity and presenting easier examples before more challenging ones [26]. In the domain of metric learning, Sanakoyeu et al. [27, 28] partition both the data and the embedding space into smaller subsets and learn separate distance metrics within each of the non-overlapping subspaces. Meanwhile, Sarkar et al. [29, 30] proposed a hierarchical sampling strategy for retrieval tasks, initially sampling negatives from the same high-level category and then selecting harder negatives from the same fine-grained category.
Figure 2: This figure shows a few examples of visually similar objects from different categories in our dataset that makes it suitable for fine-grained recognition and retrieval tasks.
@@
Pose-invariant Ranking Losses: Multi-view [31, 5, 32, 1] and pose-invariant methods [7] primarily focused on learning embeddings specific to categories, where the object-to-object variations are represented by variations in embedding vectors within the same embedding space. However, these approaches struggle to effectively distinguish between embeddings of different objects within the same category, leading to subpar performance on object-based tasks. To ameliorate this, PiRO [11] designed a dual-encoder model that explicitly decouples object and category embeddings. By training this model using pose-invariant ranking losses optimized in separate embedding spaces simultaneously, PiRO achieved improved object recognition and retrieval performance by effectively distinguishing between similar objects within the same category.
@@
In our work, we evaluate the performance of these methods on our new dataset. Additionally, we introduce a modified training approach that employs the same losses as PiRO [11], but with a curriculum learning strategy that follows a different mechanism for sampling informative examples and conducting joint training, as detailed in Sec. 4.2.1. While PiRO randomly samples a pair of objects from the same category, we additionally sample similar objects from the same and other categories. As the training progresses, we sample objects that are harder to distinguish. This approach captures more discriminative object-identity representations to distinguish between visually similar objects, leading to performance improvements on object-level tasks, as shown later in Tables III, III.
(C) Transformer-based Architectures: Hybrid architectures, combining a CNN backbone with a transformer encoder [33], have been used for learning aggregated representations from image collections comprising different items [30, 29] and multi-view 3D shapes [34, 35] for classification and retrieval tasks. PiRO [11] introduced a dual-encoder to simultaneously learn pose-invariant category and object representations. While PiRO utilizes a single-head self-attention layer for each subspace to aggregate visual features from different views, we adopt a similar architecture with multiple attention layers. This modification allows us to effectively capture invariant visual features under more complex transformations beyond just pose changes.
@@ 3.1 Dataset Design and Collection
Category and Object Selection: Objects are chosen such that they satisfy the following two conditions: (a) Objects have significant appearance changes under different state and pose changes (b) Objects are chosen such that different objects from the same category have very subtle differences for fine-grained object recognition and retrieval tasks as shown in Fig. 2. From 21 categories [^1] The categories are bags, books, bottles, bowls, clothes, cups, decorations, headphones, telephones, pillows, plants, plates, remotes, retail products, toys, ties, towels, trolley bags, tumblers, umbrellas, and vegetables., 331 objects that meet these criteria are selected. For each object, multiple images are collected based on the method specified next.
Dataset Collection: The OWSC dataset consists of images of 331 household objects captured using smartphone cameras. For each object, two sets of images are collected:
  • Transformation [^2] Note that ‘transformation’ in the context of our work refers to variations in state as well as viewpoint, pose, illumination, background, etc. split: For this split, images of each object in diverse states are captured from multiple arbitrary viewpoints. The objects are placed in multiple stable poses on different surfaces, featuring diverse backgrounds and varying lighting conditions. Detailed descriptions outlining the different transformations specific to each category are elaborated in the supplemental (ref. Table V).
  • Probe split: For this split, images of each object placed in an unseen state are captured from arbitrary viewpoints. The state of the objects, including the background and pose, differs from the images in the transformation split.
Due to differences in the physical attributes of the objects, they may undergo diverse types of transformations, resulting in a dataset with a variable number of images per object. Overall, the transformation split comprises 7900 images with an average of 24 images per object, and the probe split comprises 3428 images with approximately 10 images for each object. The transformation split is used for training, while the probe split comprising images of the same objects in an unseen state, is used for testing. Some examples are shown later in Fig. 7 and Figs. 8, 9 in the supplemental.
Dataset Annotation and Pre-processing: Each object is annotated with a category label, object label, and a text description describing the visual characteristics of the object to facilitate multi-modal learning. The images are captured from arbitrary viewpoints without recording specific pose or viewpoint information. No preprocessing is performed to preserve the natural background of the images, ensuring their suitability for in-the-wild recognition and retrieval.
3.2 Benchmarking on ObjectsWithStateChange
Single-image and Multi-image Invariant Recognition and Retrieval Tasks: We evaluate methods on eight invariant category and object-level recognition and retrieval tasks on our dataset. Single-image or multi-image category recognition. These tasks predict the category from a single image or a set of object images in an unseen state respectively. Single-image or multi-image object recognition. These tasks predict the object label from a single image or a set of object images in an unseen state respectively. Single-image or multi-image category retrieval. The goal of these tasks is to retrieve images of objects from the same category as the query object from a single image or multiple images respectively. Single-image or multi-image object retrieval. The goal of these tasks is to retrieve images of the same object-identity as the query object from a single image or multiple images respectively. Note that for the retrieval tasks, the query images (from the probe split) are in a different state as compared to the retrieved images (from the transformation split). The data splits will be publicly released at https://github.com/sarkar-rohan/ObjectsWithStateChange.
Evaluation metrics: Classification performance is measured in terms of accuracy and retrieval performance is measured using mean average precision (mAP).
@@ 3.2 Benchmarking on ObjectsWithStateChange
Figure 3: An overview of the ranking losses in [11] for the object space (top) and the category space (bottom) that we use in our work.
Figure 3: An overview of the ranking losses in [11] for the object space (top) and the category space (bottom) that we use in our work.
@@
Figure 4: An overview of the proposed training pipeline to learn invariant embeddings. Our curriculum learning strategy samples pairs of visually similar objects from the same and other categories to train the dual encoder jointly using the losses in Fig. 4.
4 Learning Invariant Embeddings
Sec. 4.1 provides a brief overview of various state-of-the-art pose-invariant methods. In our work, we evaluate their performance on our ObjectsWithStateChange (OWSC) dataset. Further, we extend the state-of-the-art method [11], as detailed in Sec. 4.2.
4.1 Background of Prior Pose-invariant Methods
The methods discussed in this section were originally designed to learn pose-invariant embeddings from multi-view images, but are adaptable to learn invariant embeddings from object image collections with diverse transformations from arbitrary views as they cluster multiple images of the same object without needing explicit pose or viewpoint information during training.
4.1.1 Single Embedding Space Methods:
Ho et al. [7] proposed three pose-invariant methods PI-CNN, PI-Proxy, and PI-TC that primarily focused on learning category-specific embeddings, where the object-to-object variations are represented by variations in embedding vectors within the same embedding space. Specifically, these methods clustered single-view embeddings close to the multi-view embedding of the same object, that is clustered close to a learned proxy embedding of its category. However, the embeddings of different objects within the same category are not explicitly separated, yielding poor performance on object-based tasks.
4.1.2 Dual Embedding Space Method:
To address this, PiRO [11] designed a dual-encoder model that explicitly decouples category and object embeddings. It comprises a shared CNN backbone, and two attention heads for learning category and object-identity representations in separate embedding spaces. The attention heads capture shared features that remain invariant across various transformations of the same object using self-attention. The encoder takes as input a set of $V$ unordered images of an object undergoing transformations and outputs $V$ single-image embeddings for each individual image, and an aggregated multi-image embedding for each embedding space.
The model is trained using three specialized margin-based losses for the object and category embedding spaces. In Fig. 4, we illustrate how the same losses in [11] can be used for learning object and category representations that are invariant to transformations and state changes. The losses are:

### v2

# supplemental

@@ Abstract
We add one more invariance — the state invariance — to the more commonly used other invariances for learning object representations for recognition and retrieval. By state invariance, we mean robust with respect to changes in the structural form of the objects, such as when an umbrella is folded, or when an item of clothing is tossed on the floor. Since humans generally have no difficulty recognizing objects despite such state changes, we are naturally faced with the question of whether it is possible to devise a neural architecture with similar abilities. To that end, we present a novel dataset, ObjectsWithStateChange, which captures state and pose variations in the object images recorded from arbitrary viewpoints. We believe that this dataset will facilitate research in fine-grained object recognition and retrieval of 3D objects that are capable of state changes. The goal of such research would be to train models capable of learning discriminative object embeddings that remain invariant to state changes while also staying invariant to transformations induced by changes in viewpoint, pose, illumination, etc. A major challenge in this regard is that instances of different objects (both within and across different categories) under various state changes may share similar visual characteristics and therefore may be close to one another in the learned embedding space, which would make it more difficult to discriminate between them. To address this, we propose a curriculum learning strategy that leverages the similarity relationships in the learned embedding space after each epoch to guide the training process. In accordance with the curriculum learning principles, during the training phase we progressively select object pairs with smaller inter-object distances in the learned embedding space, thereby gradually sampling harder-to-distinguish examples of visually similar objects, both within and across different categories. Our ablation related to the role played by curriculum learning indicates an improvement in object recognition accuracy of 7.9% and retrieval mAP of 9.2% over the state-of-the-art. We believe that this strategy enhances the model’s ability to capture discriminative features for fine-grained tasks that involve objects with state changes, leading to performance improvements on object-level tasks not only on the new dataset we present, but also on three other challenging multi-view datasets such as ModelNet40, ObjectPI, and FG3D.
@@
Deep Metric Learning, Fine-grained Object Recognition and Retrieval, Pose-invariant and State-invariant Embeddings, Curriculum Learning, Multi-modal Dataset with Multi-view Images of Objects in Different States and Text Descriptions
@@ 1 Introduction
The fact that the appearance of an object depends significantly on its pose vis-a-vis that of the camera has received much attention in the research literature. However, there exists a more confounding cause for major changes in object appearance — the change in the state of the object. Consider, for example, objects like umbrellas, books on a table that may be open or closed, furniture items that are likely to be folded up when not in use, items of clothing and linen in different forms depending on how they are stowed, etc. Fig. 1 shows some examples of such objects from our ObjectsWithStateChange dataset.
We posit that with regard to the objects we encounter in our daily lives, it is just as likely as not that you will encounter objects whose appearance depends significantly on what state we find them in. That makes important the following question: Can modern computer vision algorithms effectively recognize objects despite the changes in their state? That is, in addition to achieving pose and viewpoint invariances, is it possible to also achieve invariance with respect to state changes? We believe that the work we present is a step in the direction of answering those questions.
Figure 1: In addition to pose, many commonly occurring objects also exhibit significant changes in their appearance when their state changes. This figure displays several examples of such objects in our ObjectsWithStateChange (OWSC) dataset.
@@
Obviously, any exploration of the answers to such questions must begin with the creation of a dataset that captures not only pose and viewpoint appearance changes, but also state changes. To that end, a primary goal of this paper is the presentation of a new dataset we have created — the ObjectsWithStateChange dataset (OWSC). We believe that this dataset nicely augments the other well-known datasets — ModelNet40, ObjectPI, and FG3D — that are only meant to train and test for invariances with respect to pose and viewpoint. We believe that the OWSC dataset would facilitate creating image representations that would also be invariant to highly nonlinear shape deformations caused by state changes in addition to being invariant to changes in viewpoint, pose, illumination, etc.
To highlight the usefulness of the OWSC dataset, we first evaluate several state-of-the-art pose-invariant methods [7], [11], and conclude that these methods struggle with object recognition and retrieval tasks when objects undergo state changes. Subsequently, we propose an extension to the dual-encoder model [11] that learns discriminative object and category representations simultaneously by comparing images under different transformations of a pair of objects that are randomly sampled from the same category. However, it doesn’t allow the model to distinguish subtle differences between similar items, like half-sleeve and full-sleeve tops within the same category, or visually similar items like towels and blankets in different categories. The problem becomes even more challenging under various state changes, such as when clothing items are folded or piled. In order to accommodate such changes in object state, we extend [11] by proposing a curriculum learning strategy that uses the inter-object distances in the learned embedding space after each epoch to sample pairs of objects that have similar visual characteristics and are harder to distinguish in the presence of state changes.
Following curriculum learning principles, we progressively select object pairs with closer inter-object distances in the learned embedding space as training advances, thus gradually introducing harder-to-distinguish examples of visually similar objects within and across categories. Specifically, this strategy involves sampling images of similar objects with various state changes from the same category (by retrieving neighboring objects within the category in the learned embedding space) and from different categories (by subdividing the embedding space further as training progresses and sampling nearby objects within each partition). Although this method introduces the added complexity of finding a neighboring object for each object to form pairs, compared to simple random sampling within a category [11], it maintains sub-linear time complexity relative to the number of objects per category using fast approximate nearest neighbor techniques (see Appendix B). This approach promotes cross-category and within-category comparisons, encouraging the model to differentiate between challenging examples and enhancing its ability to capture discriminative features for fine-grained tasks while being invariant to state and pose changes. This yields significant improvements in recognition and retrieval tasks over existing methods, as shown in Sec. 5.
@@
  • We introduce a new dataset called ObjectsWithStateChange (OWSC) that incorporates state changes in object appearance that are in addition to the more commonly used transformations related to changes in pose and viewpoint. We evaluate several state-of-the-art approaches for recognition and retrieval with this dataset. The recognition and retrieval tasks are both at the object-level and at the category-level.
  • We propose a new mining strategy that is based on the principles of curriculum learning and that allows our framework to learn discriminative object-level representations more effectively that works well even in the presence of state changes for the objects. This strategy leads to performance improvements on object-level tasks not only on the proposed ObjectsWithStateChange dataset but also on three other publicly available multi-view datasets, ObjectPI, ModelNet-40, and FG3D.
Figure 2: Figs. (a)-(e) show a few examples of visually similar objects from different categories in our dataset for fine-grained recognition and retrieval tasks in the top three rows. Fig. (f) shows samples of pairs of visually similar objects from the same category under various state and pose changes captured from arbitrary views in the bottom two rows.
Figure 3: This figure shows samples from the Train and Test splits of our dataset for different objects. The state of the object as well as the background and pose are different in each split for every object. The images are captured from arbitrary viewpoints.
Figure 4: This figure shows the images in various states and poses captured from arbitrary viewpoints, category label, object label, and text description for each object of the OWSC dataset.
@@ 2 Related Work
(A) Multi-view Object Datasets: Prior work has introduced several multi-view object datasets for recognition and retrieval tasks. Datasets such as ModelNet-40 [1], and FG3D [2] comprise multi-view images of CAD models of objects. These datasets are intended for learning just the shape of objects and lack color and texture information, that we see in real-world objects. Datasets such as ETH-80 [3], W-RGBD [4], RPC [9], MIRO [12], MVP-N [10] capture real-world objects placed in a controlled environment with uniform illumination and a plain background while the ObjectPI [7] dataset captures objects in the wild with complex backgrounds. In all these datasets, multi-view data is captured for objects that are placed in particular poses, the data being recorded from a set of pre-defined viewpoints. ObjectNet [8] is a test dataset meant for evaluation of category recognition in a real-world setting with arbitrary object backgrounds, rotations, and viewpoints. Toybox [6] is a small video dataset capturing basic transformations such as rotation, translation, and zoom.
@@
(B) Curriculum Learning: As introduced in [13], it organizes learning tasks by gradually increasing the complexity of the data samples during the training phase, starting with the easiest cases and ending with those that are the most challenging [14]. As an extension of this basic idea, in the domain of metric learning, Sanakoyeu et al. [15] partition both the data and the embedding space into smaller subsets and learn separate distance metrics within each of the non-overlapping subspaces. Meanwhile, Sarkar et al. [16, 17] proposed a hierarchical sampling strategy for retrieval tasks, initially sampling negatives from the same high-level category and then selecting harder negatives from the same fine-grained category.
@@
(C) Pose-invariant Methods: Pose-invariant methods [7] focus on learning category-specific embeddings, representing the object-to-object variations within the same embedding space. However, these approaches struggle to differentiate between objects within the same category, resulting in subpar performance on object-based tasks. To address this, PiRO [11] designed a dual-encoder model to decouple object and category embeddings, achieving improved object recognition and retrieval performance by effectively distinguishing between similar objects within the same category.
@@
In our work, we evaluate the performance of these methods on our new dataset. Additionally, we train the PiRO [11] framework with a curriculum learning strategy. While PiRO randomly samples a pair of objects from the same category, we additionally sample similar objects from the same and other categories. As the training progresses, we sample objects that are harder to distinguish. This approach captures more discriminative object-identity representations to distinguish between visually similar objects, improving performance on object-level tasks (ref. Tables III(b), III(b)). Also, the dual endoder in PiRO utilizes a single-head self-attention layer for each subspace to aggregate visual features from different views, we adopt a similar architecture with multiple attention layers. This modification allows us to effectively capture invariant visual features under more complex transformations beyond just pose changes (ref. Table IV).
@@ 3.1 Dataset Design and Collection
Category and Object Selection: We chose objects for the OWSC dataset such that they satisfied the following two conditions: (a) The objects had significant appearance changes under different state and pose changes; and (b) We chose objects that we thought would be particularly challenging because the differences in their images (from different viewpoints for the same category and across categories) appeared to be subtle for fine-grained object recognition and retrieval tasks. Fig. 2 shows some examples of the images collected for such objects. The dataset contains images from the following 21 categories: bags, books, bottles, bowls, clothes, cups, decorations, headphones, telephones, pillows, plants, plates, remotes, retail products, toys, ties, towels, trolley bags, tumblers, umbrellas, and vegetables. The number of objects in these categories totals 331. We made sure that all these objects met the two conditions just stated. For each object, multiple images were collected as described below.
Dataset Collection: The OWSC dataset comprises 11328 images of 331 household objects from 21 categories, captured using smartphone cameras under various state changes. Detailed descriptions of these state variations and other transformations specific to each object category, are provided in the supplemental (ref. Table A5). Each object in a given state was placed in multiple stable poses on various surfaces, with variations in lighting conditions and backgrounds. Objects were then photographed under these variations from arbitrary viewpoints, capturing a wide range of angles and diverse perspectives.
During data collection, all images of an object in the same state captured from multiple viewpoints, including variations in pose, lighting, and background, were grouped together. The dataset was then randomly partitioned into training and test sets with a 3:1 ratio, where one-fourth of the states per object were randomly selected for the test set, and all images of the object in those states were allocated to the test set, while the images from the remaining three-fourths of the states of the same object were assigned to the training set. This ensures that the training and test sets contain images of each object in different states, with no overlap between the two partitions. The training dataset consists of 7,900 images, with approximately 24 images per object, while the test dataset contains 3,428 images, with approximately 10 images per object. This setup allows the model to be evaluated on its ability to learn invariant representations for recognizing and retrieving objects across different states and transformations. Some examples from both splits are shown in Fig. 4. The dataset is publicly released at https://github.com/sarkar-rohan/ObjectsWithStateChange.
@@
Figure 5: An overview of the ranking losses from [11] described in Sec. 4.2 for the object space (top) and the category space (bottom) that we use in our work to learn invariant embeddings.
@@
Figure 6: An overview of the proposed training pipeline to learn invariant embeddings. Our curriculum learning strategy uses inter-object distances in the currently learned embedding space to sample pairs of visually similar objects from the same and other categories to train the dual encoder jointly using the losses in Fig. 6. As shown in the left, our method switches between three sampling strategies (ref. Sec. 4.3.1), based on the epoch number $N_{e}$.
Dataset Annotation and Pre-processing: Each object is annotated with a category label, object label, and a text description describing the visual characteristics of the object to facilitate multi-modal learning, as shown in Fig. 4. The images are captured from arbitrary viewpoints without recording specific pose or viewpoint information. No preprocessing is performed to preserve the natural background of the images, ensuring their suitability for in-the-wild recognition and retrieval.
3.2 Benchmarking using ObjectsWithStateChange
Single-image and Multi-image State-Invariant Recognition and Retrieval Tasks: For evaluation using our dataset, we introduce eight state-invariant tasks focused on category- and object-level recognition and retrieval as described below:
Single-image or multi-image category recognition: Predict the category label from either a single image or a set of images.
Single-image or multi-image object recognition: Predict the object-identity label from either a single image or a set of images.
Single-image or multi-image category retrieval: Retrieve images of objects belonging to the same category as the query object, using either a single image or a set of images as the query.
Single-image or multi-image object retrieval: Retrieve images of the same object-identity as the query object, using either a single image or a set of images as the query.
Evaluation metrics: Classification and retrieval performance are measured in terms of accuracy and mean average precision (mAP).
4 Proposed Framework for Learning Invariant Embeddings
For the work reported in this paper, we adapted the PiRO framework [11] to deal with the challenges caused by state changes of objects that exhibit similar visual characteristics. First, Sec. 4.1 provides a brief overview of the network architecture and modifications we make. Next, Sec. 4.2 describes the losses in [11] and how they can be used for learning state-invariant embeddings from a collection of object images. Lastly, we elaborate on the motivation and the details of our curriculum learning strategy to mine informative samples for learning discriminative state-invariant embeddings in Sec. 4.3.
4.1 The Dual-Encoder Architecture
PiRO [11] introduced a dual-encoder model comprising a shared CNN backbone and two attention heads, each dedicated to learning category and object-identity representations in two distinct embedding spaces. In the work described in [11], using self-attention in their respective embeddings spaces, the heads captured category- and object-specific features that remained invariant to transformations like pose changes. As we will show later in this paper, the same embeddings also remain invariant to state changes when trained using curriculum learning, as described in Sec. 4.3.
In our work, we input a set of $V$ unordered images of an object corresponding to its different states and other transformations to the encoder that outputs $V$ single-image embeddings and an aggregated multi-image embedding for each embedding space. While PiRO employs a single-headed self-attention layer to aggregate features from different viewpoints, our method uses multiple attention layers, enabling it to effectively capture invariant features under intricate transformations caused by state changes, extending beyond just pose variations. This is validated in our ablation study in Table IV.
4.2 Losses
The model is trained using three margin-based losses from [11] to optimize the distances between the images of object pairs to learn state-invariant object and category embeddings, as follows: